A truck rental agency charges a daily fee plus a kilometre fee.julie was charged R460 for two days and 100 kilometres and christina was charged R1050 for three days and 400 kilometres. what is the agency's daily fee and what is the kilometres fee

Answer by addingup(3677)

2d + 100k = 460 multiply all times 3
3d + 400k = 1050 multiply all times -2 and add
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
6d + 300k = 1380
+
-6d + -800k = -2100
-----------------
0d + -500k = -720 divide by -500, and remember that -/- = +
k = 720/500 = R1.44 this is how much the agency charges per kilometer.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Using the first equation, let's find d:
2d + 100k = 460
2d + 100(1.44) = 460
2d + 144 = 460
2d = 316
d = R158 this is how much the agency charges per day
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
2d + 100k = 460
2(158) + 100(1.44) = 460 Correct (multiply and add)
.
2nd equation:
3d + 400k = 1050
3(158) + 400(1.44) = 1050 Correct